# PoseWatch: A Transformer-based Architecture for Human-centric Video Anomaly Detection Using Spatio-temporal Pose Tokenization

### v1

@@ Abstract
Video Anomaly Detection (VAD) presents a significant challenge in computer vision, particularly due to the unpredictable and infrequent nature of anomalous events, coupled with the diverse and dynamic environments in which they occur. Human-centric VAD, a specialized area within this domain, faces additional complexities, including variations in human behavior, potential biases in data, and substantial privacy concerns related to human subjects. These issues complicate the development of models that are both robust and generalizable. To address these challenges, recent advancements have focused on pose-based VAD, which leverages human pose as a high-level feature to mitigate privacy concerns, reduce appearance biases, and minimize background interference. In this paper, we introduce PoseWatch, a novel transformer-based architecture designed specifically for human-centric pose-based VAD. PoseWatch features an innovative Spatio-Temporal Pose and Relative Pose (ST-PRP) tokenization method that enhances the representation of human motion over time, which is also beneficial for broader human behavior analysis tasks. The architecture’s core, a Unified Encoder Twin Decoders (UETD) transformer, significantly improves the detection of anomalous behaviors in video data. Extensive evaluations across multiple benchmark datasets demonstrate that PoseWatch consistently outperforms existing methods, establishing a new state-of-the-art in pose-based VAD. This work not only demonstrates the efficacy of PoseWatch but also highlights the potential of integrating Natural Language Processing techniques with computer vision to advance human behavior analysis.
keywords:
Video Anomaly Detection, Human Centric, Human Pose, Privacy-Preserving, Human Behavior Analysis, Transformers
Statements and Declarations
This research is funded and supported by the United States National Science Foundation (NSF) under Award Numbers 1831795 and 2329816.
All datasets used in this manuscript are publicly available.
1 Introduction
Figure 1: A conceptual overview of PoseWatch. PoseWatch assigns higher scores to anomalous pose sequences. The final frame score is the maximum score of all individuals in the scene.
Video Anomaly Detection (VAD) is a rapidly growing area within Computer Vision that focuses on automatically identifying unusual events or behaviors in video sequences [46, 42, 1, 45]. This technology has a wide array of practical applications, including smart surveillance [43, 42], traffic monitoring [75, 69, 44], and healthcare [41]. A notable subset of VAD is human-centric anomaly detection, which specifically targets the recognition of atypical human behaviors.
@@ 1 Introduction
This paper introduces PoseWatch, a novel non-autoregressive transformer-based model with an innovative spatio-temporal tokenization approach for pose-based human-centric anomaly detection. The non-autoregressive design of PoseWatch enables the simultaneous generation of output tokens, crucial for the time-sensitive nature of anomaly detection. Figure 1 provides an abstract overview of the PoseWatch system. PoseWatch features two innovative components: the Spatio-Temporal Pose and Relative Pose (ST-PRP) tokenization and the Unified Encoder Twin Decoders (UETD) transformer core. The ST-PRP tokenization is designed to maximize the self-attention capabilities of the transformer, creating a new paradigm in pose tokenization for a range of advanced pose-based tasks. The UETD core, with its Future Target Decoder (FTD) and Current Target Decoder (CTD), processes tokens and calculates anomaly scores. This architecture combines a single unified encoder with two decoder heads, each tailored for specific operational goals.
PoseWatch, with just 0.5 million parameters, introduces a self-supervised approach to anomaly detection that sets a new standard for performance across multiple benchmark datasets, including ShanghaiTech [35], HR-ShanghaiTech [40], and the Charlotte Anomaly Dataset [13], achieving State-of-the-Art (SotA) average Area Under the Receiver Operating Characteristic Curve (AUC-ROC) score of $80.67\%$. Additionally, crucial for real-world applications, PoseWatch maintains an impressive average Equal Error Rate (EER) of $0.27$, demonstrating SotA balance between false negatives and false positives on these datasets.
@@
  • Introducing PoseWatch, the combination of a novel non-autoregressive Unified Encoder Twin Decoders (UETD) transformer and the ST-PRP tokenization, featuring Current Target Decoder (CTD) and Future Target Decoder (FTD) for self-supervised human anomaly detection.
  • Demonstrating the accuracy and generalizability of PoseWatch through comparison with not only SotA pose-based approaches but also pixel-based approaches across various benchmark datasets.
2 Related Works
@@
Figure 2: PoseWatch architecture. ST-PRP tokenization reorders and prepares input pose sequences for being fed to the UETD transformer core. The UETD transformer core consists of a unified pose transformer encoder and twin decoders for CTD and FTD. The MSE loss of both CTD and FTD branches is used to calculate the Current Score (CS) and Future Score (FS), respectively. The average of the two scores is calculated to find the Hybrid Score (HS). Please note that $a$ and $b$ are constant multipliers both set to $0.5$ for calculating the HS. Red and blue represent PoseWatch-F and PoseWatch-C data flows respectively.
3 PoseWatch
The architecture of PoseWatch, depicted in Figure 2, embodies two main components: the Spatio-Temporal Pose and Relative Pose (ST-PRP) tokenization, alongside the Unified Encoder Twin Decoders (UETD) transformer core. PoseWatch leverages a shared encoder, a Current Target Decode (CTD), and a Future Target Decoder (FTD). The CTD and FTD branches provide complementary insights by capturing distinct patterns within the input sequences. This synergy enhances the overall model’s robustness, as it minimizes the influence of individual branch errors through the aggregation of results from both branches. In the following subsections, we delve into the details of Posewatch.
3.1 ST-PRP Tokenization
@@
The transformer’s design exclusively employs inter-token self-attention, ignoring any intra-token attention mechanisms [62]. Consequently, the way attention is applied depends significantly on how the input data is tokenized. In order to utilize the full potential of the transformer self-attention module we introduce a Spatio-Temporal Pose and Relative Pose Tokenization (ST-PRP). After conducting extensive experiments with various tokenization strategies, as detailed in Section 6, we identified ST-PRP as the best-performing approach. ST-PRP tokenization, as depicted in Figure 3, employs $\beta$ tokens, each with dimensions of $k\times 2\times 2$. The initial $\beta/2$ tokens are dedicated to $x$ coordinates, and the latter half pertains to $y$ coordinates. Each token encapsulates both the absolute and relative values of either $x$ or $y$ coordinates of a pose in two adjacent frames. Spatial attention arises from the attention between keypoints and their $x$ and $y$ dimensions, whereas temporal attention is obtained by the frame number progression across tokens. Transformers, by design, do not inherently understand the order of input unless it’s explicitly provided. The ST-PRP tokenization captures spatial and temporal relationships within and across frames, but it doesn’t inherently indicate the sequence of tokens. Thus, we use the positional encoding strategy [62] to embed order into the input sequence and construct the input of the PoseWatch Unified Encoder Twin Decoders (UETD) transformer core or $TC_{i}$.
Figure 3: PoseWatch Spatio-Temporal Pose and Relative Pose (ST-PRP) Tokenization Schema. $k$ is the number of keypoints, $\beta$ is the input window size, $\Delta$ shows relative coordinates and $x(t,k)$ and $y(t,k)$ are the coordinates of $k^{th}$ keypoint in time step $t$.
Further insights into the significance and characteristics of relative pose utilization, along with a detailed analysis of various tokenization methods, through empirical evaluations, are presented in Section 6.
3.2 UETD Transformer Core
At the heart of PoseWatch is the UETD module, which processes tokens generated by the ST-PRP through a dual-branch structure. This core component is trained in a self-supervised manner, enabling effective anomaly detection. The following sections will go into the details of each of these branches and their architecture.
3.2.1 CTD Branch (PoseWatch-C):
@@
The ST-PRP Tokenization output serves as the input for PoseWatch-C, being treated as a sequential data stream. The PoseWatch CTD branch has an encoder-decoder structure. We chose our design to use a non-autoregressive strategy to take advantage of the parallelization of the transformer’s structure. In the CTD branch, we chose the target sequence of the decoder to be equal to the current sequence $TC_{i}^{t_{0}}$. Both the encoder and the decoder are chosen to have 12 heads. Considering the real-time nature of the anomaly detection task, we choose a minimal 4 number of layers with the feed-forward layer dimensions set to 64. Unlike most NLP tasks, we do not need to define a start token and end token for the sequences since the input and output sequences have fixed lengths. PoseWatch does not use masking strategies for both the input and target sequences since these sequences are always available even in the inference time. The output of the CTD branch for the input $TC_{i}^{t_{0}}$ is:
@@
3.2.2 FTD Branch (PoseWatch-F):
As illustrated in Figure 2, the PoseWatch FTD and CTD branches utilize a shared encoder. This encoder leverages an advanced understanding of pose dynamics and progression when trained for the CTD branch. Correspondingly, the FTD branch employs a decoder mirroring the architecture of the CTD, leveraging 12 heads and 4 layers. Although these decoders have identical architecture (denoted as ’Twin Decoders’), each of them is distinctively tasked. While both generate the current sequence, their inputs differ in timing with FTD’s input lagging by one sequence step compared to CTD. Following the non-autoregressive strategy, for the input sequence of $TC_{i}^{t_{0}-\beta}$, the target sequence of the decoder is the future sequence ($TC_{i}^{t_{0}}$ is considered the future sequence compared to the input of $TC_{i}^{t_{0}-\beta}$) in both the training and inferencing process. For the same reason as the CTD branch, We do not use any masking in the FTD branch either. The output of the FTD branch for the input $TC_{i}^{t_{0}-\beta}$ is:
@@ 3.2.2 FTD Branch (PoseWatch-F):
3.2.3 PoseWatch Hybrid (PoseWatch-H):
@@
In addressing the real-time demands of anomaly detection, our commitment to minimal model complexity is evident, exemplified by only choosing 4 layers for both the encoder and twin decoders. Additionally, diverging from vision transformers employed in various computer vision tasks, our strategy involves tokenized poses with reduced dimensions, resulting in PoseWatch-H only having 0.5 million parameters and an average end-to-end latency of 5.96 ms.
4 Experimental Setup
@@
4.1 Metrics
@@
4.2 Training Strategy and Hyper-partameters
@@
SHT [35] has been recorded at 24 FPS. Thus, we consider the input sequence length to be 24, equivalent to 1s. PoseWatch-C underwent training for $20$ epochs with a learning rate of $51e-$ 5. In the next step, we freeze the parameters of the trained pose encoder and train the PoseWatch-F decoder for $30$ epochs with a learning rate of $32e-$ 3.
@@
CHAD [13] is recorded at 30 FPS. Thus, we chose the input sequence length to be 1s or 30 frames. PoseWatch-C is trained for $30$ epochs with a learning rate of $32e-$ 3. In the next step, We freeze the parameters of the pose encoder and train the PoseWatch-F decoder for $30$ epochs with a learning rate of $45e-$ 4.
@@
Figure 4: AUC-ROC of PoseWatch compared with pose-based approaches on SHT [35], HR-SHT [40], and CHAD [13] datasets. PoseWatch is compared to MPED-RNN [40], GEPC [39], PoseCVAE [23], MSTA-GCN [8], MTP [48], HSTGCNN [72], STGformer [21], MoPRL [68] and STG-NF [19]. The best results of PoseWatch is demonstrated using dashed lines per each dataset. The average is only calculated for algorithms with results for all three datasets for fairness. Detailed numbers can be seen in Table 1.
5 Results
5.1 Comparison With Pose-based Approaches
Table 1 presents the performance of PoseWatch and its variants on three distinct datasets: SHT [35], HR-SHT [40], and CHAD [13]. Unfortunately, most works do not provide their code base for further examinations on the newer dataset.
@@ 5.1 Comparison With Pose-based Approaches
Figure 5: Output anomaly scores of PoseWatch-H for each frame of clip $01\_0025$ from the SHT dataset [35]. The red area on the plot indicates the ground truth anomalous frames. In this clip, the anomalous behavior is a person riding a bike on the sidewalk shown by the red rectangle.
@@
Table 1: AUC-ROC of PoseWatch compared with pose-based approaches on SHT [35], HR-SHT [40], and CHAD [13] datasets. PoseWatch is compared to SotA methods such as MPED-RNN [40], GEPC [39], PoseCVAE [23], MSTA-GCN [8], MTP [48], HSTGCNN [72], STGformer [21], MoPRL [68] and STG-NF [19]. The best is in bold and the second best is underlined.
Methods SHT HR-SHT CHAD Average
MPED-RNN 73.40 75.40 - -
GEPC 75.50 - 64.90 -
PoseCVAE 74.90 75.70 - -
MSTA-GCN 75.90 - - -
MTP 76.03 77.04 - -
HSTGCNN 81.80 83.40 - -
STGformer 82.90 86.97 - -
MoPRL 83.35 84.40 - -
STG-NF 85.90 87.40 60.60 78.63
PoseWatch-C 85.10 86.70 66.12 79.31
PoseWatch-F 83.19 83.70 66.61 77.17
PoseWatch-H 85.75 87.23 67.04 80.67
@@
Table 2: EER of PoseWatch compared with pose-based approaches on SHT [35], HR-SHT [40], and CHAD [13] datasets. PoseWatch is compared to SotA methods such as GEPC [39] and STG-NF [19]. The best is in bold and the second best is underlined.
Methods SHT HR-SHT CHAD Average
GEPC 0.31 - 0.38 -
STG-NF 0.22 0.21 0.43 0.29
PoseWatch-C 0.23 0.22 0.38 0.28
PoseWatch-F 0.25 0.25 0.38 0.29
PoseWatch-H 0.22 0.21 0.37 0.27
Based on Table 1 and Figure 4 PoseWatch-H establishes itself as the most accurate model for VAD with an average AUC-ROC that surpasses the previous state-of-the-art by $2.04\%,$ as demonstrated across three benchmark datasets. Specifically, on the CHAD dataset, PoseWatch-H outperforms the prior best model by $2.14\%$, highlighting its robustness in complex environments. On the SHT and HR-SHT datasets, PoseWatch-H secures the second-highest AUC-ROC scores of $85.75\%$ and $87.23\%,$ respectively, trailing STG-NF by only $0.15\%$ and $0.17\%$. However, STG-NF’s approach, which relies on probabilistic modeling with a fixed normal distribution, limits its generalizability, particularly on the diverse CHAD dataset, where it underperforms by $6.44\%$. This shortfall underscores PoseWatch-H’s superior ability to model diverse normal scenarios, making it the most effective model overall. Additionally, PoseWatch-C, which ranks second in average AUC-ROC, further demonstrates the effectiveness and versatility of the PoseWatch approach, solidifying its position as a new standard in pose-based VAD.
Examining the EER across all datasets, Table 2 reveals that PoseWatch-H consistently outperforms previous SotA, with an average EER of $0.27$ compared to STG-NF’s $0.29$. This trend mirrors the AUC-ROC results, further solidifying PoseWatch-H as the superior model in terms of overall performance and generalizability. Both PoseWatch-H and STG-NF achieve an EER of 0.22 on the SHT dataset. However, the CHAD dataset underscores the limitations of STG-NF, where it shows a higher EER of 0.43 compared to PoseWatch-H’s lower EER of 0.37. This difference highlights PoseWatch-H’s broader applicability and robustness across diverse datasets, making it a more versatile solution, especially for real-world applications where balancing false positive rates (usability) and false negative rates (security) is crucial.
In the comparison of PoseWatch-H with its variants, namely PoseWatch-C and PoseWatch-F, as depicted in Table 1, it is evident that PoseWatch-H consistently demonstrates superior performance when compared to its counterparts. For instance, in the SHT dataset, PoseWatch-C and PoseWatch-F register AUC-ROC scores of $85.10\%$ and $83.19\%$, respectively, whereas PoseWatch-H surpasses them with an AUC-ROC of $85.75\%$ and $87.23\%$. This enhancement in performance is attributed to the synergistic effect of integrating CTD and FTD branches. It shows that the CTD alone may miss certain anomalies, which the FTD can detect, and vice versa. This complementary relationship is not limited to AUC-ROC; similar trends are observed in EER measurements, shown in Table 2, further underscoring the robustness of PoseWatch-H compared to its individual sub-branches. Despite the lower AUC-ROC values and higher EER metrics observed in PoseWatch-C and PoseWatch-F relative to PoseWatch-H, these models still demonstrate competitive performance when benchmarked against other SotA methodologies, as evidenced in Table 2.
@@
Figure 6: Output anomaly scores of PoseWatch-H for each frame of clip $04\_093\_1$ from the CHAD dataset [13]. The red area on the plot indicates the ground truth anomalous frames. In this clip, the anomalous behavior is two people fighting shown by the red rectangle.
@@
Table 3: AUC-ROC of PoseWatch compared with pixel-based approaches on SHT [35] and HR-SHT [40] datasets. PoseWatch is compared with SotA methods such as MAAM-Net [64], GCLPT [71], BAF [17], BAF [17] + SSPCAB [47], SSMTL [16], SSMTL++v2 [5], Jigsaw-VAD [63] and NM-GAN [7]. The best is in bold and the second best is underlined.
Methods SHT HR-SHT Average

### v2

# Human-Centric Video Anomaly Detection Through Spatio-Temporal Pose Tokenization and Transformer

@@ Abstract
Video Anomaly Detection (VAD) presents a significant challenge in computer vision, particularly due to the unpredictable and infrequent nature of anomalous events, coupled with the diverse and dynamic environments in which they occur. Human-centric VAD, a specialized area within this domain, faces additional complexities, including variations in human behavior, potential biases in data, and substantial privacy concerns related to human subjects. These issues complicate the development of models that are both robust and generalizable. To address these challenges, recent advancements have focused on pose-based VAD, which leverages human pose as a high-level feature to mitigate privacy concerns, reduce appearance biases, and minimize background interference. In this paper, we introduce SPARTA, a novel transformer-based architecture designed specifically for human-centric pose-based VAD. SPARTA introduces an innovative Spatio-Temporal Pose and Relative Pose (ST-PRP) tokenization method that produces an enriched representation of human motion over time. This approach ensures that the transformer’s attention mechanism captures both spatial and temporal patterns simultaneously, rather than focusing on only one aspect. The addition of the relative pose further emphasizes subtle deviations from normal human movements. The architecture’s core, a novel Unified Encoder Twin Decoders (UETD) transformer, significantly improves the detection of anomalous behaviors in video data. Extensive evaluations across multiple benchmark datasets demonstrate that SPARTA consistently outperforms existing methods, establishing a new state-of-the-art in pose-based VAD.
Index Terms:
video anomaly detection, human-centric, pose-based anomaly detection, human behavior analysis, computer vision.
I Introduction
Video Anomaly Detection (VAD) is a rapidly growing area within Computer Vision that focuses on automatically identifying unusual events or behaviors in video sequences [1, 2, 3, 4]. This technology has a wide array of practical applications, including smart surveillance [5, 2], traffic monitoring [6, 7, 8], and healthcare [9]. An established subset of VAD is human-centric anomaly detection, which specifically targets recognizing atypical human behaviors.
Figure 1: A conceptual overview of SPARTA. SPARTA assigns higher scores to anomalous pose sequences. The final frame score is the maximum score of all individuals in the scene.
@@ I Introduction
This paper introduces SPARTA, a novel non-autoregressive transformer-based model with an innovative spatio-temporal tokenization approach for pose-based human-centric anomaly detection. The non-autoregressive design of SPARTA enables the simultaneous generation of output tokens, crucial for the time-sensitive nature of anomaly detection. fig. 1 provides an abstract overview of the SPARTA system. SPARTA features two innovative components: the Spatio-Temporal Pose and Relative Pose (ST-PRP) tokenization and the Unified Encoder Twin Decoders (UETD) transformer core. The ST-PRP tokenization is designed to maximize the self-attention capabilities of the transformer, creating a new paradigm in pose tokenization for a range of advanced pose-based tasks. The UETD core, with its Future Target Decoder (FTD) and Current Target Decoder (CTD), processes tokens and calculates anomaly scores. This architecture combines a single unified encoder with two decoder heads, each tailored for specific operational goals.
SPARTA, with just 0.5 million parameters, introduces a self-supervised approach to anomaly detection that sets a new standard for performance across multiple benchmark datasets, including ShanghaiTech [29], HR-ShanghaiTech [14], Charlotte Anomaly Dataset [30], and Northwestern Polytechnical University Campus [31], achieving State-of-the-Art (SOTA) average Area Under the Receiver Operating Characteristic Curve (AUC-ROC) score of $75.87\%$. Additionally, crucial for real-world applications, SPARTA maintains an average Equal Error Rate (EER) of $0.29$, demonstrating SOTA balance between false negatives and false positives on these datasets.
@@
  • Introducing SPARTA, the combination of a novel non-autoregressive Unified Encoder Twin Decoders (UETD) transformer and the ST-PRP tokenization, featuring Current Target Decoder (CTD) and Future Target Decoder (FTD) for self-supervised human anomaly detection.
  • Demonstrating the superior accuracy and generalizability of SPARTA through comparison with not only SOTA pose-based approaches but also pixel-based approaches across four benchmark datasets.
II Related Works
@@
Figure 2: SPARTA architecture. ST-PRP tokenization reorders and prepares input pose sequences for being fed to the UETD transformer core. The UETD transformer core consists of a unified pose transformer encoder and twin decoders for CTD and FTD. The MSE loss of both CTD and FTD branches is used to calculate the Current Score (CS) and Future Score (FS), respectively. The average of the two scores is calculated to find the Hybrid Score (HS). Please note that $a$ and $b$ are constant multipliers both set to $0.5$ for calculating the HS. Red and blue represent SPARTA-F and SPARTA-C data flows respectively.
III SPARTA
The architecture of SPARTA, depicted in Figure 2, embodies two main components: the Spatio-Temporal Pose and Relative Pose (ST-PRP) tokenization, alongside the Unified Encoder Twin Decoders (UETD) transformer core. SPARTA leverages a shared encoder, a Current Target Decode (CTD), and a Future Target Decoder (FTD). The CTD and FTD branches provide complementary insights by capturing distinct patterns within the input sequences. This synergy enhances the overall model’s robustness, as it minimizes the influence of individual branch errors through the aggregation of results from both branches. In the following subsections, we delve into the details of SPARTA.
III-A ST-PRP Tokenization
@@
The transformer’s design exclusively employs inter-token self-attention, ignoring any intra-token attention mechanisms [25]. Consequently, the way attention is applied depends significantly on how the input data is tokenized. In order to utilize the full potential of the transformer self-attention module we introduce a Spatio-Temporal Pose and Relative Pose Tokenization (ST-PRP). After conducting extensive experiments with various tokenization strategies, as detailed in Section VI, we identified ST-PRP as the best-performing approach. ST-PRP tokenization, as depicted in Figure 3, employs $\beta$ tokens, each with dimensions of $k\times 2\times 2$. The initial $\beta/2$ tokens are dedicated to $x$ coordinates, and the latter half pertains to $y$ coordinates. Each token encapsulates both the absolute and relative values of either $x$ or $y$ coordinates of a pose in two adjacent frames. Spatial attention arises from the attention between keypoints and their $x$ and $y$ dimensions, whereas temporal attention is obtained by the frame number progression across tokens. Transformers, by design, do not inherently understand the order of input unless it’s explicitly provided. The ST-PRP tokenization captures spatial and temporal relationships within and across frames, but it doesn’t inherently indicate the sequence of tokens. Thus, we use the positional encoding strategy [25] to embed order into the input sequence and construct the input of the SPARTA Unified Encoder Twin Decoders (UETD) transformer core or $TC_{i}$.
Figure 3: SPARTA Spatio-Temporal Pose and Relative Pose (ST-PRP) Tokenization Schema. $k$ is the number of keypoints, $\beta$ is the input window size, $\Delta$ shows relative coordinates and $x(t,k)$ and $y(t,k)$ are the coordinates of $k^{th}$ keypoint in time step $t$.
Further insights into the significance and characteristics of relative pose utilization, along with a detailed analysis of various tokenization methods, through empirical evaluations, are presented in Section VI.
III-B UETD Transformer Core
At the heart of SPARTA is the UETD module, which processes tokens generated by the ST-PRP through a dual-branch structure. This core component is trained in a self-supervised manner, enabling effective anomaly detection. The following sections will go into the details of each of these branches and their architecture.
III-B1 CTD Branch (SPARTA-C):
@@
The ST-PRP Tokenization output serves as the input for SPARTA-C, being treated as a sequential data stream. The SPARTA CTD branch has an encoder-decoder structure. We chose our design to use a non-autoregressive strategy to take advantage of the parallelization of the transformer’s structure. In the CTD branch, we chose the target sequence of the decoder to be equal to the current sequence $TC_{i}^{t_{0}}$. Both the encoder and the decoder are chosen to have 12 heads. Considering the real-time nature of the anomaly detection task, we choose a minimal 4 number of layers with the feed-forward layer dimensions set to 64. Unlike most NLP tasks, we do not need to define a start token and end token for the sequences since the input and output sequences have fixed lengths. SPARTA does not use masking strategies for both the input and target sequences since these sequences are always available even in the inference time. The output of the CTD branch for the input $TC_{i}^{t_{0}}$ is:
@@
III-B2 FTD Branch (SPARTA-F):
As illustrated in Figure 2, the SPARTA FTD and CTD branches utilize a shared encoder. This encoder leverages an advanced understanding of pose dynamics and progression when trained for the CTD branch. Correspondingly, the FTD branch employs a decoder mirroring the architecture of the CTD, leveraging 12 heads and 4 layers. Although these decoders have identical architecture (denoted as ’Twin Decoders’), each of them is distinctively tasked. While both generate the current sequence, their inputs differ in timing with FTD’s input lagging by one sequence step compared to CTD. Following the non-autoregressive strategy, for the input sequence of $TC_{i}^{t_{0}-\beta}$, the target sequence of the decoder is the future sequence ($TC_{i}^{t_{0}}$ is considered the future sequence compared to the input of $TC_{i}^{t_{0}-\beta}$) in both the training and inferencing process. For the same reason as the CTD branch, We do not use any masking in the FTD branch either. The output of the FTD branch for the input $TC_{i}^{t_{0}-\beta}$ is:
@@ III-B2 FTD Branch (SPARTA-F):
III-B3 SPARTA Hybrid (SPARTA-H):
@@
In addressing the real-time demands of anomaly detection, our commitment to minimal model complexity is evident, exemplified by only choosing 4 layers for both the encoder and twin decoders. Additionally, diverging from vision transformers employed in various computer vision tasks, our strategy involves tokenized poses with reduced dimensions, resulting in SPARTA-H only having 0.5 million parameters and an average end-to-end latency of 5.96 ms.
IV Experimental Setup
IV-A Datasets
Early datasets such as CUHK Avenue [72], Subway [73], and UCSD [74] have been foundational for VAD research. However, their limited scale has prompted the research community to adopt more complex and comprehensive recent datasets, which aligns with our focus on utilizing these modern datasets for experimentation.
@@ IV-A Datasets
Northwestern Polytechnical University Campus (NWPUC) [31] dataset encompasses 43 scenes, 28 classes of anomalous events, and 16 hours of video footage, making it a large datasets in its field. A notable feature of the NWPUC dataset is its inclusion of scene-dependent anomalies, where an event may be considered normal in one scene but abnormal in another. This dataset includes anomaly classes that are not specific to humans, presenting a disadvantage for pose-based anomaly detection methods. However, as a new and comprehensive benchmark, it was chosen for its broad applicability. This limitation affects all pose-based methods, ensuring fair comparisons.
IV-B Metrics
@@
IV-C Training Strategy and Hyper-partameters
@@
SHT [29] has been recorded at 24 FPS. Thus, we consider the input sequence length to be 24, equivalent to 1s. SPARTA-C underwent training for $20$ epochs with a learning rate of $51e-$ 5. In the next step, we freeze the parameters of the trained pose encoder and train the SPARTA-F decoder for $30$ epochs with a learning rate of $32e-$ 3.
@@
CHAD [30] is recorded at 30 FPS. Thus, we chose the input sequence length to be 1s or 30 frames. SPARTA-C is trained for $30$ epochs with a learning rate of $32e-$ 3. In the next step, We freeze the parameters of the pose encoder and train the SPARTA-F decoder for $30$ epochs with a learning rate of $45e-$ 4.
NWPUC [31] is recorded at 25 FPS. Thus, we chose the input sequence length to be 24 to be close to 1s as possible. SPARTA-C is trained for $30$ epochs with a learning rate of $35e-$ 3. In the next step, We freeze the parameters of the pose encoder and train the SPARTA-F decoder for $30$ epochs with a learning rate of $31e-$ 3.
@@
V Results
V-A Comparison With Pose-based Approaches
@@
Figure 4: Output anomaly scores of SPARTA-H for each frame of clip $01\_0025$ from the SHT dataset [29]. The red area on the plot indicates the ground truth anomalous frames. In this clip, the anomalous behavior is a person riding a bike on the sidewalk, shown by the red rectangle.
@@
TABLE I: AUC-ROC of SPARTA compared with pose-based approaches on SHT [29], HR-SHT [14], CHAD [30], and NWPUC [31] datasets. SPARTA is compared to SOTA methods such as MPED-RNN [14], GEPC [12], PoseCVAE [20], MSTA-GCN [78], MTP [21], HSTGCNN [22], STGformer [79], MoPRL [23] and STG-NF [13]. The best is in bold and the second best is underlined.
Methods SHT HR-SHT CHAD NWPUC Average
MPED-RNN 73.40 75.40 - - -
GEPC 75.50 - 64.90 62.04 -
PoseCVAE 74.90 75.70 - - -
MSTA-GCN 75.90 - - - -
MTP 76.03 77.04 - - -
HSTGCNN 81.80 83.40 - - -
STGformer 82.90 86.97 - - -
MoPRL 83.35 84.40 66.81 61.92 74.12
STG-NF 85.90 87.40 60.60 62.56 74.11
SPARTA-C 85.10 86.70 66.12 62.69 75.15
SPARTA-F 83.19 83.70 66.61 62.29 73.94
SPARTA-H 85.75 87.23 67.04 63.48 75.87
@@
TABLE II: EER of SPARTA compared with pose-based approaches on SHT [29], HR-SHT [14], CHAD [30], and NWPUC [31] datasets. SPARTA is compared to SOTA methods such as GEPC [12], STG-NF [13], and MoPRL [23]. The best is in bold and the second best is underlined.
Methods SHT HR-SHT CHAD NWPUC Average
GEPC 0.31 - 0.38 0.41 -
STG-NF 0.22 0.21 0.43 0.40 0.31
MoPRL 0.24 0.23 0.38 0.40 0.31
SPARTA-C 0.23 0.22 0.38 0.41 0.31
SPARTA-F 0.25 0.25 0.38 0.40 0.32
SPARTA-H 0.22 0.21 0.37 0.39 0.29
SPARTA-H achieves the highest average AUC-ROC, surpassing the previous SOTA by $1.75\%$ across four benchmark datasets (Table I). It outperforms the prior best model by $0.23\%$ and $0.92\%$ on CHAD and NWPUC, respectively, and ranks second on SHT and HR-SHT, trailing STG-NF by only $0.15\%$ and $0.17\%$, whose reliance on a fixed normal distribution limits its generalizability, particularly on other diverse datasets. These results highlight SPARTA-H’s robustness and versatility, making it the most effective model overall.
Examining the EER across all datasets (Table II), SPARTA-H achieves the lowest average EER of $0.29$, outperforming previous SOTA models. This aligns with the AUC-ROC results, reaffirming SPARTA-H’s superior performance and generalizability. SPARTA-H matches STG-NF on SHT and HR-SHT but demonstrates better robustness on CHAD and NWPUC with lower EER values, highlighting its versatility and effectiveness for real-world applications that require a balance between usability and security.
As shown in Table I, SPARTA-H outperforms its variants, SPARTA-C and SPARTA-F, consistently across datasets. For instance, on the SHT dataset, SPARTA-H achieves an AUC-ROC of $85.75\%$, compared to SPARTA-C’s $85.10\%$ and SPARTA-F’s $83.19\%$, demonstrating the effectiveness of its dual-branch design. The complementary synergy of CTD and FTD branches allows SPARTA-H to detect anomalies that might be missed by either branch alone. This advantage extends to EER (Table II), further underscoring the robustness and importance of the dual-branch architecture, which enhances performance and generalizability.
@@
Figure 5: Output anomaly scores of SPARTA-H for each frame of clip $04\_093\_1$ from the CHAD dataset [30]. The red area on the plot indicates the ground truth anomalous frames. In this clip, the anomalous behavior is two people fighting shown by the red rectangle.
@@
TABLE III: AUC-ROC of SPARTA compared with pixel-based approaches on SHT [29] and NWPUC [31] datasets. SPARTA is compared with SOTA methods such as MNAD [80], OG-Net [81], MemAE [82], MAAM-Net [15], MPN [83], LLSH [84], GCLPT [16], BAF [85], BAF [85] + SSPCAB [18], SSMTL [17], SSMTL++v2 [63], Jigsaw-VAD [19] and NM-GAN [86]. The best is in bold and the second best is underlined.
Methods SHT NWPUC Average
MNAD 70.50 62.50 66.50
OG-Net - 62.50 -
MemAE 71.20 61.90 66.55